OpenStudy (pinknabastak):

simplify: answer has to be arithmetic, simplified, no "i" in denominator, and in simplest form. 1) \[\frac{ 1 }{ (3i-1)(\frac{ 7 }{ 2 } i +3) }\] 2) \[\frac{ 7 }{ 3i+6 } \times \frac{ 4 }{ 9i(16+i ) }\]

OpenStudy (pinknabastak):

@agent0smith @UsukiDoll

OpenStudy (anonymous):

You can use conjugate to get the answer. 1) \[\frac{ (3i+1)(\frac{ 7 }{ 2 }i-3) }{ (3i-1)(\frac{ 7 }{ 2 }i+3)(3i+1)(\frac{ 7 }{ 2 }i-3) }\] \[= \frac{ -\frac{ 21 }{ 2 } - 3 - 5.5i }{ (9i^2 -1)(\frac{ 49 }{ 4 }i^2-9) }\] \[= \frac{ -13.5 - 5.5i }{ (-10)(-21.25) }\] \[= \frac{ -13.5 - 5.5i }{ 212.5}\] \[= \frac{ -27 - 11i }{ 425}\] You can do the same way with the second equation

OpenStudy (pinknabastak):

O.O how did you get all that?

OpenStudy (pinknabastak):

help...

hartnn (hartnn):

know what conjugate is ?

OpenStudy (pinknabastak):

no

hartnn (hartnn):

conjugate of \(a-bi = a+bi \) to find conjugate, you just flip the sign of imaginary term

OpenStudy (pinknabastak):

uh...

hartnn (hartnn):

and if you have a fraction like, 1/(a+bi) in order to make the denominator as real, we multiply and divide by the conjugate of denominator because \((a+bi)(a-bi)= a^2 +b^2\)

hartnn (hartnn):

you didn't get what conjugate is ?

OpenStudy (pinknabastak):

ok right there i got lost

OpenStudy (pinknabastak):

no i get that

OpenStudy (pinknabastak):

just not this last one

hartnn (hartnn):

know what 'i' is ?

OpenStudy (pinknabastak):

imaginary

hartnn (hartnn):

yes, i is an imaginary unit, \(\large i = \sqrt {-1}\) so can you tell me whats i^2 = ... ?

OpenStudy (pinknabastak):

\[\sqrt{-1}^{2}\]

hartnn (hartnn):

which equals ?

OpenStudy (pinknabastak):

my calc says nan

hartnn (hartnn):

lol lets not use calculator :) NAN means Not a Number (what your calc meant was, its not a REAL number, and we already know i is an imaginary number)

OpenStudy (pinknabastak):

ok so....

hartnn (hartnn):

hint : \(\large (\sqrt a)^2 = a \\ \large (\sqrt {-1})^2 =...? \)

OpenStudy (pinknabastak):

-1

OpenStudy (pinknabastak):

ok

hartnn (hartnn):

did some correction. YES, i^2 = -1 remember this. now lets go to your question, before doing 1) lets just try 1/(3i-1)  note that i multiplied numerator and denominator by conjugate of 3i-1 which is -3i+1